you are building a rectangular flower garden. you want the area of the garden to be 39 ft^2. you want the length of the garden to be 10 ft longer than its width. what dimensions of the garden should you use ?
Answer by Boreal(15235)

width=x
length=x+10
x(x+10)=39 ft^2
x^2+10x=39
x^2+10x-39=0
(x+13)(x-3)=0
x=-13,3
Only x=3 feet, width, makes sense.
Length is 13 feet
Product is 39 ft^2